OpenStudy (anonymous):

Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of f[g(5)]?

OpenStudy (anonymous):

first evaluate g(5) = 5^2 -1 =25-1 =24 then evaluate f(g(5)) = f(24) = 3*24-2 =70

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

can u make it simpler

OpenStudy (anonymous):

they are asking you to evaluate f(x) =3x-2 but they want x to be the vaule which is equal to g(5)

OpenStudy (anonymous):

soo 3x(5) -2 right

OpenStudy (anonymous):

which is 13

OpenStudy (anonymous):

x= g(5) = 5^2 -1 =24 f(24) = 3*24- 2 = 70

OpenStudy (anonymous):

the answer should be 70

OpenStudy (anonymous):

soo then wat do u do

OpenStudy (anonymous):

f(g(5)) is not a multiplication is a comp function

OpenStudy (anonymous):

can u show me step by step

OpenStudy (anonymous):

how did u get 70

OpenStudy (anonymous):

sorry I am having prob with the internet

OpenStudy (anonymous):

Let see first what is given ; f(x) = 3x-2 , g(x) = x^2 -1 they are asking to find f(g(5)) =? this means that you need to first evaluate g(x) =x^2-1 when x =5 that is g(5) = 5^2 -1 =25-1=24 because g(5) =24 this values is the one that need to plug in f(x) that is f(g(5) = f(24) = 3*24-2 =72-2=70